OpenStudy (anonymous):

x/ sin112 = 165 / sin 18.......... What is this simplified?

OpenStudy (e.mccormick):

\[\frac{x}{\sin 112}=\frac{165}{\sin 18}\] OK, so law of sines. If you multiply through by \(\sin 112\) you will have x alone and can solve for x at that point.

OpenStudy (anonymous):

So x=589.3 ?

OpenStudy (e.mccormick):

I got a different answer. hmmm.... \[\frac{x}{\sin 112}=\frac{165}{\sin 18}\implies x=\frac{165\sin 112}{\sin 18}\]

OpenStudy (anonymous):

Wow i set it up wrong... weekend brain .. sorry so its 495.1

OpenStudy (e.mccormick):

\(sin(112) \approx .9272\) and \(sin(18)\approx .3090\)

OpenStudy (e.mccormick):

Yah, that is what I got.

OpenStudy (anonymous):

thanks!!!

OpenStudy (e.mccormick):

I have made my share of mistakes and will continue to do so... we all do. The goal is to just minimize them!

OpenStudy (e.mccormick):

\[\downarrow mistakes = \uparrow grades\]

OpenStudy (anonymous):

do you think you could help me with another law of sines problem. I'm on the law of sines unit in math and i just don't get it but then when someone explains it to me its better.

OpenStudy (e.mccormick):

Well, the law of sines is all about ratios. I bet the book went through a proof, but it comes down to a basic ratio which is closely related to a fraction. Do you get that part?

OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 38°, a = 19, c = 10

OpenStudy (e.mccormick):

OK. So start with a triangle. Sketch it, fill in what you know. For example: